Assume the following, c = .15, rr = .10, and e = .2. The multiplier in this case is:   

3.67

3.33

3.7

2.56
Answer #1

Answer

Formula :

Multiplier = (c + 1)/(c + rr + e)

where c = currency to deposit ratio = 0.15, rr = required reserve ratio = 0.10 and e = excess reserve ratio = 0.2

=> Multiplier = (c + 1)/(c + rr + e) = (0.15 + 1)/(0.15 + 0.10 + 0.2) = 2.56

Hence, the correct answer is (d) 2.56